OpenStudy (anonymous):

P=12x+20y in. find an expression for the area of a square with the given perimeter

OpenStudy (dumbcow):

since we are talking about a square then all sides are equal length A = s^2 P=4s 4s = 12x+20y s = 3x+5y s^2 = (3x+5y)^2 = 9x2+15xy+25y^2 A = 9x2+15xy+25y^2